# Pruning for Protection: Increasing Jailbreak Resistance in Aligned LLMs Without Fine-Tuning

### v2

@@ 4.2 Qualitative Comparison
We also qualitatively looked into responses generated by all the models. Figure 2 Shows one example response from the base model and the pruned models. We did not find a noticeable degradation in the quality of the responses generated by the pruned models. We noticed that all model outputs, including the base model outputs, were less informative and less malicious for the most complex jailbreaking prompts, such as GAME and TOMNJERRY, and more informative and malicious for easier prompts, such as CHARACTER and KEVIN.
@@ 4.3 Benchmarking Evaluation
Table 1 summarizes our findings for Llama-2 Chat model. The corresponding benchmark results for Vicuna 1.3 and Mistral Instruct v0.2 are provided in Appendix C. We observe that generally, the pruned models perform competitively with and sometimes even surpass the base model. Since we do not observe significant degradation in the pruned models’ capabilities for reasoning, context length, or language modeling, the increased jailbreaking resistance from pruning LLaMA-2 and Vicuna cannot be attributed to a decrease in understanding of the task.
@@ 5.1 Pruning sharpens attention patterns
We inspect attention patterns and qualitatively observe that pruned models have sharper attention. Vig & Belinkov (2019) found that the entropy of attention patterns correlates with high-level semantic behavior: across various model depths, both the entropy of the attention patterns and their role in understanding sequence semantics evolve. Following this work, we calculate the entropy of attention patterns and average it over all prompts in our harmful tasks dataset, across layers and attention heads. In Figure 4, we illustrate the difference in average entropies between base and pruned models, noting that this reduction in average entropy reaches a plateau at a 20% prune percentage.
@@ 5.2 Sharper attention focuses on malicious tokens
Building on the observation that pruned models exhibit sharper attention patterns, we further analyze the distribution of attention across tokens. We measure the extent to which non-malicious ‘jailbreak’ tokens distract the model from focusing on malicious tokens. Following Vig & Belinkov (2019), we introduce a metric to capture the proportion of total attention that malicious tokens direct towards fellow malicious tokens. For every tokenized prompt $x$ in our dataset $\mathcal{X}$, we perform one forward pass and collect attention patterns $\alpha^{(l,h)}$ for every layer $l$ and attention head $h$. For a tokenized prompt $x$, we denote the set of indices originating from the original malicious task $\mathcal{T}_{x}$, while the remaining indices correspond to the different jailbreak pretexts. We introduce:
@@ 5.3 Perplexity Analysis
We now adopt an orthogonal approach to analyze, at a higher level of abstraction, how pruning influences language modeling capabilities. Our findings indicate that moderate pruning does not significantly impact language modeling performance on WikiText or the effective context length in AltQA tasks. However, this observation may not necessarily extrapolate to artificial constructs such as jailbreak templates. Indeed, it might even be preferable to have language models that do not overfit to such out-of-distribution prompts.
We approach this by investigating the perplexity assigned by both base and sparse models, to both the original malicious tasks and the prompts constructed using jailbreak templates. Note that model responses are not included in the following perplexity calculations.
In particular, for each original malicious task, we examine its perplexity before and after the application of jailbreak templates. For the latter, we report the perplexities associated with jailbreak attempts by calculating the average over the values obtained from the 10 jailbreak methods we examined.
@@
We present our results in Figure 6. Sparse models consistently assign higher perplexity scores to jailbreak constructs than base models, when both assign similar perplexities to the corresponding original malicious tasks. This increased perplexity indicates that sparse models are more sensitive to deviations from the expected distribution of language, suggesting that WANDA acts as an effective regularizer. As demonstrated in Table 1, WANDA does not incur performance penalties when modeling in-distribution language passages. In contrast, it successfully detects out-of-distribution constructs.
@@ 6 Effects of WANDA Pruning on Linear Models with Correlated Input Features
Consider a set of inputs $X^{(d\times n)}$ with correlated features, true coefficients $w^{(1\times d)}$, and target $Y^{(1\times n)}$. Assume an i.i.d. white noise $\epsilon^{(1\times n)}\sim\mathcal{N}(0,\sigma^{2})$, leading to $Y=wX+\epsilon$. We take the ordinary least square (OLS) estimate of $w$ as $w^{\mathrm{OLS}}=((XX^{T})^{-1}XY^{T})^{T}$.
Let $X=(x^{(1)},\ldots,x^{(n)})$ and $Y=(y^{(1)},\ldots,y^{(n)})$, where $x^{(1)},\ldots,x^{(n)}$ are the input data points and $y^{(1)},\ldots,y^{(n)}$ are the corresponding outputs.
@@
To generate a training dataset, first we sample a vector $\boldsymbol{\mathrm{x}}^{(1\times n)}\sim\mathcal{N}(0,1)$ and add perturbations $\delta^{(d\times n)}\sim\mathcal{N}(0,\alpha^{2})$ to it, resulting in $X_{\mathrm{train}}^{(d\times n)}=\boldsymbol{\mathrm{x}}^{(1\times n)}+\delta^{(d\times n)}$. The $\alpha$ controls the level of correlation in the input features. A low $\alpha$ indicates a high correlation among the input features and vice versa. After that, we sample $\epsilon^{(1\times n)}\sim\mathcal{N}(0,\sigma^{2})$ and create $Y_{\mathrm{train}}^{(1\times n)}=w^{(1\times d)}X_{\mathrm{train}}^{(d\times n)}+\epsilon^{(1\times n)}$. We sample another $\boldsymbol{\mathrm{x}}^{(1\times n)}$ and repeat the process for the test dataset.
Next, for each trial, we obtain $w^{\mathrm{OLS}}$ using the training samples, apply WANDA to prune 30% of the weights of $w^{\mathrm{OLS}}$, and then compare the MSE loss of the unpruned and the pruned estimators on the test dataset.
@@ 7 Conclusion
In this work, we explored the effects of pruning on the jailbreaking resistance of large language models. By applying Wanda pruning at varying levels of sparsity to LLaMA-2-7B-Chat, Vicuna 1.3, and Mistral Instruct v0.2 models, we obtained an assortment of compressed models with which to compare the susceptibilities of jailbreaking. We further curated a dataset of 225 malicious tasks and 2250 jailbreaking prompts, with which we evaluated our base and compressed models. Our results show that if the unpruned model is sufficiently safety trained, then safety improves at lower sparsities of pruning, but then a reversal in the trend when pruned more aggressively. This suggests the possibility of using a carefully selected amount of pruning to aid in the deployment of safe LLMs.
@@
8 Impact Statement
This paper presents work whose goal is to introduce a novel method for improving the resistance of LLMs to producing harmful content. The societal consequences of our work may, however, extend beyond our original intentions. Open-sourcing a dataset of jailbreaking prompts and malicious tasks could potentially enable the generation of more dangerous content. We acknowledge this risk, but with the rapid growth in usage of LLMs and the need for safety as well, we hope that our work and dataset will enable and promote further research in this area.
@@ Appendix C Details about the Benchmarks
  • MMLU: Massive Multitask Language Understanding (MMLU) is a comprehensive benchmark encompassing a wide range of subjects and domains, from humanities to natural sciences, intended to evaluate an AI model’s broad understanding and reasoning capabilities across diverse topics. (Hendrycks et al., 2021)
  • TruthfulQA: TruthfulQA is designed to assess the ability of language models to provide truthful and factual answers. (Lin et al., 2022) It consists of questions that are intentionally misleading or prone to the elicitation of falsehoods, testing the model’s resistance to propagating inaccuracies.
  • Winograde: The Winograde Schema Challenge is a natural language understanding test focusing on coreference resolution, where the task is to resolve ambiguity in sentences that require understanding the relationships between different entities. (Sakaguchi et al., 2019)
  • GSM8k: Grade School Math 8k (GSM8k) is a benchmark consisting of grade-school level math problems, designed to evaluate an AI’s capability in understanding and solving basic arithmetic and mathematical reasoning questions. (Cobbe et al., 2021)
@@
WikiText(Merity et al., 2016) 5.938 5.938 5.941 5.944
Table 4: Vicuna 1.3 performance on 7 key benchmarks. Scores excluding perplexity are presented in %. Evaluation on the AltQA(Pal et al., 2023) 2k-token benchmark is omitted due to exceeding the maximum sequence length of the model.

### v3

@@ 4.2 Qualitative Comparison
We also qualitatively analyzed the responses generated by all the models. Figure 2 presents an example response from the base model alongside the pruned model’s. We did not observe a significant degradation in response quality for the pruned models. Interestingly, across all models—including the base models—the outputs were less informative and less malicious for the more complex jailbreaking prompts, such as GAME and TOMNJERRY, while they tended to be more informative and malicious for simpler prompts like CHARACTER and KEVIN.
@@ 4.3 Benchmarking Evaluation
Table 1 summarizes our findings for the LLaMA-2 Chat model. The corresponding benchmark results for Vicuna 1.3 and Mistral Instruct v0.2 are provided in Appendix C. Overall, we find that the pruned models perform competitively with, and sometimes even outperform, the base model. Since we did not observe significant degradation in reasoning, context handling, or language modeling capabilities, the increased jailbreaking resistance observed in the pruned LLaMA-2 and Vicuna models cannot be attributed to a reduction in task understanding.
5 Automatic prompt generation attacks
5.1 GCG
We evaluate how pruning enhances safety robustness against automatic prompt generation attacks. Zou et al. (2023) introduced GCG, a greedy gradient-based search method for generating adversarial prompt suffixes. They evaluated this attack across multiple scenarios, including attacking a single white-box model to generate harmful outputs and transferring adversarial suffixes to black-box models. In our study, we focus on the single-model setup and examine how pruning defends against the attack’s ability to induce harmful behaviors.
Model Success Fail $p$ value
Llama2 4 6 N/A
Llama2 10% pruned 5 5 0.65
Llama2 20% pruned 4 6 1.00
Llama2 30% pruned 0 10 0.03
Llama2 40% pruned 4 6 1.00
Table 2: Pruning at 30% sparsity enhances model robustness against GCG-generated adversarial prompts in the single-model setup.
Using the LLaMA-2 model and its variants pruned at 10%, 20%, 30%, and 40% target sparsity, we reevaluated the models and present our results in Table 2. Due to computational constraints, we evaluated only the first 10 examples from the AdvBench harmful behavior dataset. We manually labeled all completions and allowed GCG to run for 500 steps for each target behavior. To assess whether pruning led to statistically significant safety improvements, we computed $p$-values to determine if the differences in successful attack rates between models could be attributed to chance, assuming the successes follow a Bernoulli distribution. Our analysis revealed that pruning at 30% target sparsity induces statistically significant safety improvements. We believe that the safety enhancement peaks at a higher sparsity level than in manual jailbreak scenarios because GCG attacks are more efficient, requiring stronger regularization to maintain the models’ safety filters.
5.2 Advbench within our jailbreaks
We also evaluated the refusal rates of LLaMA-2 models on jailbroken prompts derived from AdvBench. Our findings indicate that our dataset is more effective at triggering malicious responses than AdvBench itself. Table 3 presents the number of refusals out of 5,720 malicious requests.
Model base 10% 20% 30% sparse
Refusals 5699 5704 5706 5695
Table 3: Refusal counts of LLaMA-2 models against AdvBench harmful behaviors embedded within our 10 jailbreak templates. Safety improvements peak at 20% sparsity, similar to our findings with the previously introduced malicious task dataset.
@@ 6.1 Pruning sharpens attention patterns
We inspect attention patterns and qualitatively observe that pruned models have sharper attention. Vig and Belinkov (2019) found that the entropy of attention patterns correlates with high-level semantic behavior: across various model depths, both the entropy of the attention patterns and their role in understanding sequence semantics evolve. Following this work, we calculate the entropy of attention patterns and average it over all prompts in our harmful tasks dataset, across layers and attention heads. In Figure 4, we illustrate the difference in average entropies between base and pruned models, noting that this reduction in average entropy reaches a plateau at a 20% prune percentage.
@@ 6.2 Sharper attention focuses on malicious tokens
Building on the observation that pruned models exhibit sharper attention patterns, we further analyze the distribution of attention across tokens. We measure the extent to which non-malicious ‘jailbreak’ tokens distract the model from focusing on malicious tokens. Following Vig and Belinkov (2019), we introduce a metric to capture the proportion of total attention that malicious tokens direct towards fellow malicious tokens. For every tokenized prompt $x$ in our dataset $\mathcal{X}$, we perform one forward pass and collect attention patterns $\alpha^{(l,h)}$ for every layer $l$ and attention head $h$. For a tokenized prompt $x$, we denote the set of indices originating from the original malicious task $\mathcal{T}_{x}$, while the remaining indices correspond to the different jailbreak pretexts. We introduce:
@@ 6.3 Perplexity Analysis
We now adopt an orthogonal approach to analyze, at a higher level of abstraction, how pruning influences language modeling capabilities. Our findings indicate that moderate pruning does not significantly impact language modeling performance on WikiText. However, this observation may not necessarily extrapolate to artificial constructs such as jailbreak templates. Indeed, it might even be preferable to have language models that do not overfit to such out-of-distribution prompts.
We approach this by investigating the perplexity assigned by both base and sparse models, to both the original malicious tasks and the prompts constructed using jailbreak templates. Note that model responses are not included in the following perplexity calculations. For each original malicious task, we examine its perplexity before and after the application of jailbreak templates. For the latter, we report the perplexities associated with jailbreak attempts by calculating the average over the values obtained from the 10 jailbreak methods we examined.
@@
We present our results in Figure 6 for the 20% sparse Llama2 model. The sparse model consistently assign higher perplexity scores to jailbreak constructs than base models, when both assign similar perplexities to the corresponding original malicious tasks. This increased perplexity indicates that sparse models are more sensitive to deviations from the expected distribution of language, suggesting that WANDA acts as an effective regularizer. As demonstrated in Table 1, WANDA does not incur performance penalties when modeling in-distribution language passages. In contrast, it successfully detects out-of-distribution constructs.
@@ 7 Effects of WANDA Pruning on Linear Models with Correlated Input Features
Consider a set of inputs $X^{(d\times n)}$ with correlated features, true coefficients $w^{(1\times d)}$, and target $Y^{(1\times n)}$. Assume an i.i.d. white noise $\epsilon^{(1\times n)}\sim\mathcal{N}(0,\sigma^{2})$, leading to $Y=wX+\epsilon$. We take the ordinary least square (OLS) estimate of $w$ as $w^{\mathrm{OLS}}=((XX^{T})^{-1}XY^{T})^{T}$. Let $X=(x^{(1)},\ldots,x^{(n)})$ and $Y=(y^{(1)},\ldots,y^{(n)})$, where $x^{(1)},\ldots,x^{(n)}$ are the input data points and $y^{(1)},\ldots,y^{(n)}$ are the corresponding outputs.
@@
To generate a training dataset, first we sample a vector $\boldsymbol{\mathrm{x}}^{(1\times n)}\sim\mathcal{N}(0,1)$ and add perturbations $\delta^{(d\times n)}\sim\mathcal{N}(0,\alpha^{2})$ to it, resulting in $X_{\mathrm{train}}^{(d\times n)}=\boldsymbol{\mathrm{x}}^{(1\times n)}+\delta^{(d\times n)}$. The $\alpha$ controls the level of correlation in the input features. A low $\alpha$ indicates a high correlation among the input features and vice versa. After that, we sample $\epsilon^{(1\times n)}\sim\mathcal{N}(0,\sigma^{2})$ and create $Y_{\mathrm{train}}^{(1\times n)}=w^{(1\times d)}X_{\mathrm{train}}^{(d\times n)}+\epsilon^{(1\times n)}$. We sample another $\boldsymbol{\mathrm{x}}^{(1\times n)}$ and repeat the process for the test dataset. Next, for each trial, we obtain $w^{\mathrm{OLS}}$ using the training samples, apply WANDA to prune 30% of the weights of $w^{\mathrm{OLS}}$, and then compare the MSE loss of the unpruned and the pruned estimators on the test dataset.
@@ 8 Conclusion
In this work, we explored the effects of pruning on the jailbreaking resistance of large language models. By applying WANDA pruning at varying levels of sparsity to LLaMA-2-7B-Chat, Vicuna 1.3, and Mistral Instruct v0.2 models, we obtained an assortment of compressed models. We further curated a dataset of 225 malicious tasks and 2250 jailbreaking prompts, with which we evaluated our base and compressed models. Our results show that if the unpruned model is sufficiently safety trained, then safety improves at lower sparsities of pruning, but then a reversal in the trend when pruned more aggressively. This suggests the possibility of using a carefully selected amount of pruning to aid in the deployment of safe LLMs.
@@ Appendix C Details about the Benchmarks
  • MMLU: Massive Multitask Language Understanding (MMLU) is a comprehensive benchmark encompassing a wide range of subjects and domains, from humanities to natural sciences, intended to evaluate an AI model’s broad understanding and reasoning capabilities across diverse topics. Hendrycks et al. (2021)
  • TruthfulQA: TruthfulQA is designed to assess the ability of language models to provide truthful and factual answers. Lin et al. (2022) It consists of questions that are intentionally misleading or prone to the elicitation of falsehoods, testing the model’s resistance to propagating inaccuracies.
  • Winograde: The Winograde Schema Challenge is a natural language understanding test focusing on coreference resolution, where the task is to resolve ambiguity in sentences that require understanding the relationships between different entities. Sakaguchi et al. (2019)
  • GSM8k: Grade School Math 8k (GSM8k) is a benchmark consisting of grade-school level math problems, designed to evaluate an AI’s capability in understanding and solving basic arithmetic and mathematical reasoning questions. Cobbe et al. (2021)
@@
WikiTextMerity et al. (2016) 5.938 5.938 5.941 5.944
Table 6: Vicuna 1.3 performance on 7 key benchmarks. Scores excluding perplexity are presented in %. Evaluation on the AltQAPal et al. (2023) 2k-token benchmark is omitted due to exceeding the maximum sequence length of the model.